The rest mass of an electron is 9.1 x


10^-31 kg. What is the relativistic mass if its velocity is 55 % of the speed


of light?

1
Expert's answer
2022-02-04T18:00:35-0500

"m=\frac{m_0}{\sqrt{1-\frac{v^2}{c^2}}}=10.9\cdot 10^{-31}~kg."